OpenStudy (anonymous):

what adds up to 3 and multiplies to -30?

OpenStudy (tkhunny):

A couple of different things. Which one would you like and how do you propose we find it?

OpenStudy (anonymous):

the same 2 numbers that do both things

OpenStudy (anonymous):

1

OpenStudy (tkhunny):

Okay, that doesn't quite mean enough. Do you mean this? Take the number 124. This "adds up" to 1+2+4 = 7 This "multiplies" to 1*2*4 = 8 I hope you don't mean that, since we might have a hard time with the negative result.

OpenStudy (anonymous):

Are you sure its not adds up to 13? I can't figure anything else out. 15 and -2 if so...

OpenStudy (anonymous):

no I mean ex. -14n-51 ----- (n-17)(n+3)=0 -------- -17*3=-51 and -17+3=-14

OpenStudy (anonymous):

then that makes our lives difficult lol and did u mean +17 +3=14 in your example?

OpenStudy (anonymous):

oh see nvm

OpenStudy (anonymous):

its easy what equation adds up to 3 but also multiplies to -30?

OpenStudy (anonymous):

then u do it :p

OpenStudy (anonymous):

I cant figure it out duhh

OpenStudy (anonymous):

i thought it was easy :p

OpenStudy (anonymous):

ur smart im not that much

OpenStudy (anonymous):

aww dont say that :( give me a sec :)

OpenStudy (anonymous):

just do the dang problem lol :b

OpenStudy (anonymous):

try bumping your question to the top again :p

OpenStudy (anonymous):

wait, are you trying to factorize x^2 +3x -30=0?

OpenStudy (anonymous):

cause that cannt be factorized